The population of a city increases at a rate proportional to the present number. It has an initial population of 50000 that increases by 15% in 10 years. What will be the population in 30

years?


1
Expert's answer
2022-04-15T05:23:12-0400

Define "x(t)" - population of the city at the moment "t" (in years).

"\frac{dx}{dt}=mx" - reflect the fact that population of the city increases at a rate proportional to the present number. "m" is proportional coefficient. Let’s solve this equation:

"\frac{dx}{x}=mdt"

"\ln |x|=mt+\ln C"

"x=Ce^{mt}" , where "C=const" .

"x(0)=C=50000" ;

"x(10)=50000+0.15\cdot50000=57500"

"x(10)=50000e^{10m}=57500"

"e^{10m}=\frac{57500}{50000}=1.15"

"10m=\ln{1.15}"

"m=\frac{\ln{1.15}}{10}\approx0.014"

"x(t)=50000e^{\frac{\ln{1.15}}{10}t}"

Population in 30 years:

"x(30)=50000e^{\frac{\ln{1.15}}{10}\cdot30}=50000\cdot(1.15)^3=" "76043.75\approx76044"